OpenStudy (anonymous):

If one angle of a set of supplementary angles measures 85°, then the other angle measures

OpenStudy (unklerhaukus):

the sum of supplementary angles is 180°

OpenStudy (theviper):

Let the other angle be 'x' then according to @UnkleRhaukus :-\[\LARGE{x+85^o=180^o}\]\[\Large{x=180^o-85^o=??}\]